OpenStudy (anonymous):

The number of State Parks in California is 36 more than one third the number of State Parks in Florida. There are 85 State Parks in California. Which of the following equations would you use to find the number of State Parks in Florida? Note: f represents the number of state parks in Florida. 3(f − 36) = 85 f + 36 = 85

OpenStudy (hba):

he number of State Parks in California is 36 more than one third the number of State Parks in Florida. 85 is 36 more than one third the number of State Parks in Florida. 85 = 36 + one third the number of State Parks in Florida 85 = 36 + f/3 this is the equation I would use. Of course the equation can be simplified and solved, and you are not showing us your list of choices. If I were to continue this problem, I would subtract 36 from both sides and get another equation, 49 = f/3 f = 147

OpenStudy (anonymous):

the first one ...

OpenStudy (anonymous):

Thanks guys. (: